OpenStudy (lolipop_drum312):

Enter the solution to the equation in the box. 1/3(x−7)=−1/6(x−10) x =

alones (alones):

So you have to simplify?

OpenStudy (lolipop_drum312):

idk

OpenStudy (calculusxy):

\[\frac{1}{3} \times \frac{(x-7)}{1} = -\frac{1}{6} \times \frac{(x-10)}{1} \]

OpenStudy (calculusxy):

\[\frac{(x-7)}{3} = -\frac{(x-10)}{6}\]

OpenStudy (calculusxy):

\[6(x-7) = 3 \times -(x-10) \rightarrow 6x - 42 = 3(-x+10)\] \[6x - 42 = 3(-x+10) \rightarrow 6x - 42 = -3x + 30\] \[6x - 42 = -3x + 30 \rightarrow 9x - 42 = 30 \rightarrow 9x = 72 \] \[\frac{9x}{9} = \frac{72}{9} \rightarrow \large \bf \color{green}{x=8}\]

OpenStudy (i_am_legend):

The answer is x=8 hope this helps

alones (alones):

Dooh it'd be obvious..'cause @calculusxy said the answer already.. @i_am_legend and you're not suppose to say DA's...at least explain like calculuxy did.

OpenStudy (i_am_legend):

I didn't see his answer dummy, and Idc I can say answers if I want everybody fine with it anyway now mind your own work

alones (alones):

Bruh, ig you didn't read the rules.. and no need to say "dummy" guess you're blind than.